OpenStudy (anonymous):

Choose all answers that are correct. 2. (1 pt) Which equations show that the set of whole numbers is closed under addition? Choose all answers that are correct. A. 0 + (–1) = –1 B. 1 + 1 = 2 C. 2 + 0 = 2 D. 2 + (–1) = 1

OpenStudy (anonymous):

please help will fa and medal :)

OpenStudy (daniellelovee):

why is there a 2 in the question (does it means that you have to find for 2) because they are all correct

OpenStudy (daniellelovee):

and were is the set of numbers?

OpenStudy (anonymous):

its one question but you have to enter in all that are correct, Thank you for the help :)

OpenStudy (anonymous):

@Daniellelovee so they are all correct?

OpenStudy (danjs):

0 + (-1) = -1 -1 is in the set of {0,-1}

OpenStudy (danjs):

1 + 1 = 2 2 is not in the set of {1,1}

OpenStudy (danjs):

It is closed under addition, if when you add any two elements in the set, you result in another element that is always in the set.

OpenStudy (danjs):

2 + 0 = 2 2 is in the set of {2,0}

OpenStudy (anonymous):

A & D?

OpenStudy (danjs):

2 + -1 = 1 1 is not in the set of {2,-1}

OpenStudy (funnyguy2):

not allowed to help you on a test

OpenStudy (anonymous):

@DanJS Just A? Are you sure there are no more, it can be more than 1 answer.

OpenStudy (anonymous):

oh sorry :/ i thought this you could on here @funnyguy2